OpenStudy (anonymous):

5a^3/a^7 How do I solve this? I'm getting 5/a^-7

OpenStudy (anonymous):

sorry I messed up the question, its actually 5a^3/a^7

OpenStudy (anonymous):

5/a^-4

OpenStudy (anonymous):

I know thats the answer but how do I get there? What am I doing wrong because I'm getting 5/a^-7

OpenStudy (anonymous):

errr on the answer key it says the answer is 5/a^4

OpenStudy (ghazi):

\[\frac{ 5a^3 }{ a^7 }=\frac{ 5 }{ a^{7-3} }=\frac{ 5 }{ a^4 }=5a^{-4}\]

OpenStudy (anonymous):

Hmmm I see you're subtracting the denominator exponent by the numerator exponent. In my book examples it has the opposite where for the example question x^5/x^2 they subtract the numerator exponent by the denominator exponent to get x^3. Is there some mathematical rule I'm missing?

OpenStudy (ghazi):

nope, actually if you should subtract from the bigger to the smaller power , there is no set of rule for this

OpenStudy (ghazi):

lets say this is an example \[\frac{ a^4 }{ a^2 }=a^{4-2}=a^2\] in other case you can take another example \[\frac{ a^3 }{ a^5 }=\frac{ 1 }{ a^{5-3} }=\frac{ 1 }{ a^2 }\]

OpenStudy (ghazi):

this way you can get it done :D

OpenStudy (anonymous):

Oh alright, thanks for explaining it to me. I get it now.

OpenStudy (ghazi):

you are welcome :)